OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (-1,5) m= -1

OpenStudy (anonymous):

y -y(1) = m(x-x(1)) , where (x1, y1) = (-1,5) and m = (-1) Therefore, y-5 = (-1)(x+1) y-5 = -x-1 y-5+5 = -x -1+5 y = -x+4, where m=(-1) and b=4